Shift Summary Day, August 5, 1999

Yee Bob Hsiung (hsiung@fnal.gov)
Thu, 5 Aug 1999 15:47:38 -0500 (CDT)

Shift Summary for day shift, 5/Aug/99

On Shift: Yee Bob Hsiung, John Shield

Hours of beam delivered at all intensities: 8.5
Hours of beam delivered at nominal intensity: 8.5
Hours of E832 data (special runs not included): 4hrs 45 min
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: 3hrs 45 min

Numerous drt_fresh_starts and daq problems after changing new tapes
in the morning (see below).

========================================================================
Runs ended this shift: 14109
Run 14109 - 6.4 M - BEAM_832 (tape full)
Run 14117 - 13 spills - BEAM_832 (aborted and overwritten maybe due to one
bad tape in ktev1 drv3)
Run 14118 - in progress - BEAM_832

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Smooth running until 8:20am tape changing then all hell breaks loose
on daq problems till noon resulting in the loss of 3hr 45 min of
beam time. Smooth running in the afternoon part of shift.
The daq problems started with ktev4a disk 95% full (no warning) which
prevented us to start new run (pink box error - ocpr_da: error starting
run xxxx), followed by hangup of kumquat crates which requires hardware
reset, and then followed by timeouts on gateway and DAMP processes.
Finally we had one successful drt_fresh_start for run 14117, but
got aborted by run control after 13 spills data due to a possible
bad tape in ktev1 drive tps2d2 for tape QKS219. Replaced it with
a new tape and get run 14118 going smoothly around noon.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14109 7:10 YBH, JS took over the smooth run.
labeled 10 new QKB081-090 tapes.
~95-100 7:50 pipeline error, rest pipeline the kfrend.
ktev1a disk full warning, clean up files.
14109 8:20 Run ended due to tapes full.
changing tapes.

8:30-11:00 Problems with daq to start new run as described
above. Paged and phoned Julie, Ed and Viv.
Julie came in to rescue.
drt_fresh_start failed numerous time.

14117 11:02 Finally we struggled through and get a new
run going.
14117 13 11:18 Run aborted by itself.
Maybe a bad tape QKS219 on ktev1 drv3.
After Paging and consulting with Sydney, we
replaced it with a new tape and
overwrite the data.


14118 1 11:55 new run on QK*219-221 and QKB077.
13:05 ktevana error, but kfrend did not work.
power cycled laser repeater downstairs
but got crate 53 busy, Jason H unplugged
and re-plugged the DYC command cable to
force init from kfrend. It works.

13:30 NickS gave tour to U.Chicago summer students.

104 14:13 ktevana error, kfrend.

========================================================================
Follow up Notes on Previous Shift

o Hot wire in kumquat 3Y-6 is being monitored and is not causing any
problems w/ trigger inefficiencies. There is no need to page the
expert unless frowny faces in "Kumquat Readout Errors" plot turn to
dead faces.

o minor Etot red botton in ktrig window.
========================================================================
Items to Follow up on Next Shift

o At the next tape change, please do a muon run and change magnet
polarity.
o continue taking E832 data at 6.8E12.